News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Monday 13 September 1999 7.00am EST

During the last 24 hours the on-board load (P13) continued operation and included a coordinated observation of Capella with the Hubble Space Telescope, EUVE, BeppoSAX and the Very Large Array (VLA).

The procedure to re-insert the Low Energy Grating was completed successfully and initial assessment is that a software timing correction is needed to ensure the Mechanism Control Electronics (MCE) is turned off after retraction. The MCE was turned off manually following a grating retraction after the Capella observation.

Assessment and analysis of the ACIS front-side energy resolution degradation continued. The next test with ACIS is to perform a bakeout of the instrument and raise the focal plane to 25 degrees C. This procedure was started at 4.30AM 9/13 and is planned to take 24 hours. The load was stopped in preparation for the procedure and the remaining HRC calibration observations re-planned for a later mission schedule. The bakeout sequence is proceeding nominally.





Return to Status Reports